FACTORISING

RICHARD P. BRENT

Abstract

A popular article describing the factorisation of the Fermat numbers F_{8} and F_{11}.

Comments

Only an Abstract is given here. The paper appeared as [4]. For the original papers, see [1, 2, 3].

References

1. R. P. Brent and J. M. Pollard, “Factorization of the eighth Fermat number”, Mathematics of Computation 36 (1981), 627–630. MR 83h:10014. rpb061.

2. R. P. Brent, “Factorization of the eleventh Fermat number (preliminary report)”, AMS Abstracts 10 (1989), 89T–11–73. rpb113a.

3. R. P. Brent, “Parallel algorithms for integer factorisation”, Number Theory and Cryptography (edited by J. H. Loxton), London Mathematical Society Lecture Note Series 154, Cambridge University Press, 1990, 26–37. MR 91h:11148. rpb115.

4. R. P. Brent, “Factorising”, Australian Mathematical Society Gazette 16, 5 (October 1989), 154–155. rpb118.

Computer Sciences Laboratory, Australian National University, Canberra E-mail address: rpb@cslab.anu.edu.au

1991 Mathematics Subject Classification. Primary 11-04; Secondary 11A51, 11Y05, 11Y11, 14H52. Key words and phrases. Fermat number, factorization. Copyright \begin{matrix}c \\ ◯ \end{matrix} 1989, Australian Mathematical Society. Comments \begin{matrix}c \\ ◯ \end{matrix} 1993, R. P. Brent. rpb118a typeset using AMS-L^{A}TEX